OpenStudy (anonymous):

Multiply. Express your answer in lowest terms. 3/8 X 2/7 =

OpenStudy (anonymous):

3*2=6 8*7=56

OpenStudy (anonymous):

2 goes into both

OpenStudy (anonymous):

confused on the other steps

OpenStudy (anonymous):

are these fractions or are you deviding?

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

fractions

OpenStudy (anonymous):

You only use terms for fractions

OpenStudy (anonymous):

3/8 X 2/7 =3/28

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw

OpenStudy (secret-ninja):

@Zoe9639 Is correct. :) Good job both of you. ^-^

OpenStudy (anonymous):

:D